OpenStudy (anonymous):

Maya and Sally were the two finalists in a singing competition. The person with the most votes from the audience was chosen as the winner. Sally received 25% of the total votes and lost by 62,500 votes. No person in the audience was allowed to vote more than once. The equation below represents this situation, where y is the total number of votes cast and x is the number of votes Maya received. x - 62,500 = 0.25y If a total of 125,000 people voted, how many votes did Maya receive? Do not enter comma for placeholder values. @sodukogirl13

OpenStudy (anonymous):

@sodukogirl13 can you help

OpenStudy (anonymous):

x-62500=0.25*125000 x-62500=31250 x=93750

OpenStudy (anonymous):

A store offers a 15% discount on all items in the store during a sale. All store employees get an additional 10% employee discount. In the equation below, C represents the cost of an item for an employee including tax, and x represents the original price of the item. C = 0.765x + 0.06(0.765x) Chris is an employee at the store. He plans to buy a coat during the sale. The price of the coat before the discount is $100. How much will Chris pay for the coat during the sale? Remember, do not include the $ symbol.

OpenStudy (anonymous):

C=0.765*100+0.06(0.765*100) C=76.5+0.06*76.5 C=76.5+4.59 C=81.09

OpenStudy (anonymous):

The sum of three consecutive integers is 267. What is the largest integer? @sodukogirl13

OpenStudy (anonymous):

267/3=89 88+89+90=267 90